The coin is tossed m times. Find the probability p that the head appears at least k times and at most l times


1
Expert's answer
2022-03-31T13:44:10-0400

There are 2m possible results. "C^k_m" is the number of results with k heads. Probability of k heads is "\frac{C^k_m}{2^m}" .

So probability from k till l heads is

"P=\sum^l_{i=k}\frac{C^i_m}{2^m}"